# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v4

@@ Appendix B Additional related work
The mentioned inversion might seem a technicality but it actually constitutes a fundamental point in the study of IGB. One observable that characterizes the phenomenon is the fraction of points classified as class $c$, and denoted by $f_{c}\left(\mathcal{W}\right)$, (this quantity will be formally introduced shortly) in fact does not respect the self-averaging property [Mézard et al., 1987, Dotsenko, 1994]. [^7] In a system defined over an ensemble of realizations (in our case, each realization is a different weight initialization), a self-averaging quantity is one that can be equivalently calculated either by averaging over the whole ensemble or on a single, sufficiently large, realization; in such a situation, a single huge system is adequate to represent the entire ensemble. In other words, even in the infinite dataset/network size limit, the value of $f_{0}\left(\mathcal{W}\right)$ obtained on a given realization of the network initialization, $\mathcal{W}$, differs from that computed from the average over the ensemble of initializations. Self-averaging is often exploited (e.g. in dynamical mean field theory [Sompolinsky et al., 1988], employed also in the context of mean-field theories of deep learning [Schoenholz et al., 2016, Poole et al., 2016, Xiao et al., 2018, Yang and Schoenholz, 2017, Yang et al., 2019]), as it can lead to a simplification of the analysis. For the phenomena related to IGB we cannot exploit self-averaging.
@@ C.1.2 CLT extension
For a set of independent r.v.s $\{X_{i}\}_{i=1}^{n}$, if Lindeberg’s condition holds for all positive $\epsilon$, then, defining $S_{n}=X_{1}+\dots+X_{n}$ and $Z_{n}=\frac{S_{n}-\sum_{i}\mu_{i}}{s_{n}}$, we have $p_{Z_{n}}\left({z}\right)\xrightarrow{n\rightarrow\infty}\mathcal{N}\left(z;0,1\right)\,.$ (25)
@@
Let us consider a set of independent zero-mean r.v.s $\{X_{i}\}_{i=1}^{n}$. [^8] In general if $\mathbb{E}\left({X_{i}}\right)\neq 0$ we can define a new set of variables $\{\left(X_{i}-\mathbb{E}\left({X_{i}}\right)\right)\}$. If and only if, for every fixed $\epsilon>0$, the following conditions are satisfied: Concentration: $\sum_{i=1}^{n}\mathbb{P}\left(|X_{i}|\geq\epsilon\right)\xrightarrow{n\rightarrow\infty}0\;\;\;\;\forall\epsilon\in\mathbb{R}^{+}\,$, (26) Mean Normalization: $\sum_{i}\left(\int_{|x|<\epsilon}x\,p_{X_{i}}\left({x}\right)dx\right)\xrightarrow{n\rightarrow\infty}0\,$, (27) Variance Normalization: $s_{n}^{2}=\sum_{i}\left(\int_{|x|<\epsilon}x^{2}\,p_{X_{i}}\left({x}\right)dx-\left(\int_{|x|<\epsilon}x\,p_{X_{i}}\left({x}\right)dx\right)^{2}\right)\xrightarrow{n\rightarrow\infty}\sigma^{2}\,$, (28) the distributions of the sum $\sum_{i}X_{i}$ will converge to $\mathcal{N}\left(0,\sigma^{2}\right)$.
@@
Let us consider a set $\{X_{i}\}_{i=1}^{n}$ of independent, zero-mean, [^9] In general if $\mathbb{E}\left({X_{i}}\right)\neq 0$ we can define a new set of variables $\{\left(X_{i}-\mathbb{E}\left({X_{i}}\right)\right)\}$. r.v.s satisfying, for some constants $\tilde{\sigma}^{2}\in\mathbb{R}^{+}$, the following conditions Variance convergence: $\sum_{i=1}^{n}\mathbb{E}\left({X_{i}^{2}-\mathbb{E}\left({X_{i}}\right)^{2}}\right)\xrightarrow{n\rightarrow\infty}n\tilde{\sigma}^{2}\,$, (29) Fast decreasing tails: $\lim_{x\rightarrow\pm\infty}\,p_{X_{i}}\left({x}\right)=\mathcal{O}\left(\frac{1}{x^{4}}\right),\;\;\forall i\,.$ (30) Let us define the new set $\{\tilde{X}_{i}\}_{i=1}^{n}$ such that $\tilde{X}_{i}\equiv\frac{X_{i}}{c\sqrt{n}}\,$, (31) where $c>0$ is a constant. We now proof that the set $\{\tilde{X}_{i}\}_{i=1}^{n}$ satisfy (26), (27), (28) leading to the convergence in distribution of $\tilde{S}_{n}=\sum_{i}\tilde{X}_{i}$ to $\mathcal{N}\left(0,\sigma^{2}\right)$.
@@
$\begin{split}&\sum_{i=1}^{n}\mathbb{P}\left(|\tilde{X}_{i}|\geq\epsilon\right)=\sum_{i=1}^{n}\left(\int_{-\infty}^{-\epsilon}p_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}+\int_{\epsilon}^{\infty}p_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}\right)=\\ &\sum_{i=1}^{n}\left(\int_{-\infty}^{-c\sqrt{n}\epsilon}p_{X_{i}}\left({x}\right)dx+\int_{c\sqrt{n}\epsilon}^{\infty}p_{X_{i}}\left({x}\right)dx\right)\stackrel{{\scriptstyle n\gg 1}}{{=}}\\ &\sum_{i=1}^{n}\left(\int_{-\infty}^{-c\sqrt{n}\epsilon}\mathcal{O}\left(\frac{1}{x^{4}}\right)dx+\int_{c\sqrt{n}\epsilon}^{\infty}\mathcal{O}\left(\frac{1}{x^{4}}\right)dx\right)=\sum_{i=1}^{n}\left(\mathcal{O}\left(\int_{-\infty}^{-c\sqrt{n}\epsilon}\left|\frac{1}{x^{4}}\right|dx\right)+\mathcal{O}\left(\int_{c\sqrt{n}\epsilon}^{\infty}\left|\frac{1}{x^{4}}\right|dx\right)\right)=\\ &\sum_{i=1}^{n}\mathcal{O}\left(\frac{1}{n^{\frac{3}{2}}}\right)=\mathcal{O}\left(\frac{1}{n^{\frac{1}{2}}}\right)\xrightarrow{n\rightarrow\infty}0.\end{split}$ (32)
@@
$\begin{split}&\sum_{i}^{n}\left(\int_{|\tilde{x}|<\epsilon}\tilde{x}\,p_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}\right)\stackrel{{\scriptstyle n\gg 1}}{{=}}\sum_{i}^{n}\underbrace{\mathbb{E}\left({\tilde{X}_{i}}\right)}_{=0}-\sum_{i}^{n}\left(\mathcal{O}\left(\int_{-\infty}^{-c\sqrt{n}\epsilon}\frac{1}{c\sqrt{n}}\left|\frac{1}{x^{3}}\right|dx\right)+\mathcal{O}\left(\int_{c\sqrt{n}\epsilon}^{\infty}\frac{1}{c\sqrt{n}}\left|\frac{1}{x^{3}}\right|dx\right)\right)=\\ &=\mathcal{O}\left(\frac{1}{n^{\frac{1}{2}}}\right)\xrightarrow{n\rightarrow\infty}0\,.\end{split}$ (33)
@@
$\sum_{i}\left(\int_{|\tilde{x}|<\epsilon}\tilde{x}^{2}\,p_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}-\left(\int_{|\tilde{x}|<\epsilon}\tilde{x}\,p_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}\right)^{2}\right)\stackrel{{\scriptstyle n\gg 1}}{{=}}$ (34)
@@
Then we apply these results to analyze the distribution $p_{h^{(l+1)}_{i}}^{(\chi)}\left({x}\right)$; fixed $\mathcal{W}$, in fact, $h^{(l+1)}_{i}$ can be expressed as a combination of r.v.s. We will demonstrate how the nodes $\left\{h^{(l+1)}_{i}\right\}_{i}$ may not be identically distributed for $l\geq 1$.
@@ C.2.1 Combination of random variables
To examine the changes in node distribution from one layer to the next, we analyze the distribution of $h^{(l+1)}_{i}=\sum_{j}w^{(l)}_{ij}g^{(l)}_{j}$, with a fixed set $\mathcal{W}$. This leads to a linear combination of r.v.s. It is straightforward to show that the hypotheses of Thm. C.3 are satisfied in this context. We then have
$p_{h^{(l+1)}_{i}}^{(\chi)}\left({x}\right)=\mathcal{N}\left(x;\sum_{j}w^{(l)}_{ij}\left\langle{g^{(l)}_{j}}\right\rangle,\sum_{j}\left(w^{(l)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({g^{(l)}_{j}}\right)\right)$. (39)
Our goal is to understand how $p_{h^{(l+1)}_{i}}^{(\chi)}\left({x}\right)$ changes across the set $\left\{h^{(l+1)}_{i}\right\}_{i=1}^{N_{l+1}}$. Since we are dealing with a Gaussian distribution, it suffices to examine how its mean and variance vary over different nodes $\left\{h^{(l+1)}_{i}\right\}_{i=1}^{N_{l+1}}$, i.e., using different sets of weights $\left\{w^{(l)}_{i\;\boldsymbol{\cdot}}\right\}_{i=1}^{N_{l+1}}$, with $w^{(l)}_{i\;\boldsymbol{\cdot}}\equiv\left\{w^{(l)}_{ij}\right\}_{j=1}^{N_{l}}$.
Consider the mean $\left\langle{h^{(l+1)}_{i}}\right\rangle=\sum_{j}w^{(l)}_{ij}\left\langle{g^{(l)}_{j}}\right\rangle$. Now, the set $\left\{\left\langle{g^{(l)}_{j}}\right\rangle\right\}_{j=1}^{N_{l}}$ acts as a fixed constant set, [^10] This is the same set of elements $\forall i\in[1,\dots,N_{l+1}]$. while the set $\left\{w^{(l)}_{i\;\boldsymbol{\cdot}}\right\}_{j=1}^{N_{l}}$ changes with the node index $i$. Again, Thm. C.3 is easily seen to be applicable. We get
@@
Assuming $\left\langle{g^{(l)}_{j}}\right\rangle^{2}$ follows the CLT, [^11] The law of large numbers would suffice here as we are not studying convergence to the mean value. we have
$\frac{1}{N_{l}}\sum_{j=1}^{N_{l}}\left\langle{g^{(l)}_{j}}\right\rangle^{2}\xrightarrow{N_{l}\rightarrow\infty}\mathbb{E}_{w^{(l-1)}_{j\boldsymbol{\cdot}}}\left({\left\langle{g^{(l)}_{j}}\right\rangle^{2}}\right)\,$, (41)
@@
$p_{\left\langle{h^{(l+1)}_{i}}\right\rangle}^{(w^{(l)}_{i\boldsymbol{\cdot}})}\left({x}\right)\xrightarrow{N_{l}\rightarrow\infty}\mathcal{N}\left(x;0,\sigma_{w}^{2}\mathbb{E}_{w^{(l-1)}_{j\boldsymbol{\cdot}}}\left({\left\langle{g^{(l)}_{j}}\right\rangle^{2}}\right)\right)\,.$ (42)
Regarding the variance in Eq. (39), it will also be a r.v. dependent on the set $\{w^{(l)}_{ij}\}_{j=1}^{N_{l}}$. Following a similar approach to the one used for the sum of Gaussian r.v.s discussed before, we can calculate its mean and variance to show that asymptotically $\sum_{j}\left(w^{(l)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({g^{(l)}_{j}}\right)$ converges to a deterministic value equal for all $N_{l+1}$ nodes.
To make the discussion more concrete we will discuss now, more in detail the specific case of $l=3$ This will be used as an integral part of deriving $p_{f_{0}}\left({x}\right)$ for deep architectures (see App. G.3.1). In App. G.3.1, we demonstrate that
@@
Our aim is to apply the analysis described above to discern the differences between the set of distributions $\left\{p_{h^{(3)}_{i}}^{(\chi)}\left({x}\right)\right\}$. Variations in these distributions are induced by differences in the corresponding sets of weight vectors $\{w^{(2)}_{i\;\boldsymbol{\cdot}}\}$. The sets $\left\{\left\langle{g^{(2)}_{j}}\right\rangle\right\}$ and $\left\{\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)\right\}$, however, remain consistent across each node $h^{(3)}_{i}$. Therefore, we treat the latter sets as fixed random coefficients, while the variations are attributed to $\{w^{(2)}_{i\;\boldsymbol{\cdot}}\}$. Let us focus on the variance:
$\textrm{Var}_{\chi}\left({h^{(3)}_{i}}\right)=\sum_{j=1}^{N_{2}}\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)\,$. (44)
To understand how much the random variable $\textrm{Var}_{\chi}\left({h^{(3)}_{i}}\right)$ varies, we calculate its mean and variance:
$\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\sum_{j=1}^{N_{2}}\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)}\right)=\sigma_{w}^{2}\frac{1}{N_{2}}\sum_{j=1}^{N_{2}}\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)\,\xrightarrow{N_{2}\rightarrow\infty}\sigma_{w}^{2}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)}\right)$ (45)
@@
$\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)}\right)=\int_{\mathbb{R}}\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)(x)\;\;\mathcal{N}\left(x;\left\langle{h^{(2)}_{j}}\right\rangle,\sigma_{w}^{2}\left\langle{g^{(1)}}\right\rangle^{2}\right)dx\,$, (46)
where with $\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)(x)$ we underlined the dependence of $\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)$ from the r.v. $\left\langle{h^{(2)}_{j}}\right\rangle$. Note that in the integrand of Eq. (46) there is no scaling dependence with respect to $N_{2}$; this means that
$\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)}\right)=\mathcal{O}(1)\Longrightarrow\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({h^{(3)}_{i}}\right)}\right)=\mathcal{O}(1).$ (47)
To evaluate the order of magnitude of the fluctuations, we recall that, given a fixed coefficient $\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)$,
$\textrm{Var}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)}\right)=\textrm{Var}_{\chi}\left({g^{(2)}_{j}}\right)^{2}\textrm{Var}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left(w^{(2)}_{ij}\right)^{2}}\right)\,.$ (48)

### v5

@@ Appendix B Additional related work
The mentioned inversion might seem a technicality but it actually constitutes a fundamental point in the study of IGB. One observable that characterizes the phenomenon is the fraction of points classified as class $c$, and denoted by $g_{c}\left(\mathcal{W}\right)$, (this quantity will be formally introduced shortly) in fact does not respect the self-averaging property [Mézard et al., 1987, Dotsenko, 1994]. [^7] In a system defined over an ensemble of realizations (in our case, each realization is a different weight initialization), a self-averaging quantity is one that can be equivalently calculated either by averaging over the whole ensemble or on a single, sufficiently large, realization; in such a situation, a single huge system is adequate to represent the entire ensemble. In other words, even in the infinite dataset/network size limit, the value of $G_{0}\left(\mathcal{W}\right)$ obtained on a given realization of the network initialization, $\mathcal{W}$, differs from that computed from the average over the ensemble of initializations. Self-averaging is often exploited (e.g. in dynamical mean field theory [Sompolinsky et al., 1988], employed also in the context of mean-field theories of deep learning [Schoenholz et al., 2016, Poole et al., 2016, Xiao et al., 2018, Yang and Schoenholz, 2017, Yang et al., 2019]), as it can lead to a simplification of the analysis. For the phenomena related to IGB we cannot exploit self-averaging.
@@ C.1.2 CLT extension
For a set of independent r.v.s $\{X_{i}\}_{i=1}^{n}$, if Lindeberg’s condition holds for all positive $\epsilon$, then, defining $S_{n}=X_{1}+\dots+X_{n}$ and $Z_{n}=\frac{S_{n}-\sum_{i}\mu_{i}}{s_{n}}$, we have $f_{Z_{n}}\left({z}\right)\xrightarrow{n\rightarrow\infty}\mathcal{N}\left(z;0,1\right)\,.$ (25)
@@
Let us consider a set of independent zero-mean r.v.s $\{X_{i}\}_{i=1}^{n}$. [^8] In general if $\mathbb{E}\left({X_{i}}\right)\neq 0$ we can define a new set of variables $\{\left(X_{i}-\mathbb{E}\left({X_{i}}\right)\right)\}$. If and only if, for every fixed $\epsilon>0$, the following conditions are satisfied: Concentration: $\sum_{i=1}^{n}\mathbb{P}\left(|X_{i}|\geq\epsilon\right)\xrightarrow{n\rightarrow\infty}0\;\;\;\;\forall\epsilon\in\mathbb{R}^{+}\,$, (26) Mean Normalization: $\sum_{i}\left(\int_{|x|<\epsilon}x\,f_{X_{i}}\left({x}\right)dx\right)\xrightarrow{n\rightarrow\infty}0\,$, (27) Variance Normalization: $s_{n}^{2}=\sum_{i}\left(\int_{|x|<\epsilon}x^{2}\,f_{X_{i}}\left({x}\right)dx-\left(\int_{|x|<\epsilon}x\,f_{X_{i}}\left({x}\right)dx\right)^{2}\right)\xrightarrow{n\rightarrow\infty}\sigma^{2}\,$, (28) the distributions of the sum $\sum_{i}X_{i}$ will converge to $\mathcal{N}\left(0,\sigma^{2}\right)$.
@@
Let us consider a set $\{X_{i}\}_{i=1}^{n}$ of independent, zero-mean, [^9] In general if $\mathbb{E}\left({X_{i}}\right)\neq 0$ we can define a new set of variables $\{\left(X_{i}-\mathbb{E}\left({X_{i}}\right)\right)\}$. r.v.s satisfying, for some constants $\tilde{\sigma}^{2}\in\mathbb{R}^{+}$, the following conditions Variance convergence: $\sum_{i=1}^{n}\mathbb{E}\left({X_{i}^{2}-\mathbb{E}\left({X_{i}}\right)^{2}}\right)\xrightarrow{n\rightarrow\infty}n\tilde{\sigma}^{2}\,$, (29) Fast decreasing tails: $\lim_{x\rightarrow\pm\infty}\,f_{X_{i}}\left({x}\right)=\mathcal{O}\left(\frac{1}{x^{4}}\right),\;\;\forall i\,.$ (30) Let us define the new set $\{\tilde{X}_{i}\}_{i=1}^{n}$ such that $\tilde{X}_{i}\equiv\frac{X_{i}}{c\sqrt{n}}\,$, (31) where $c>0$ is a constant. We now proof that the set $\{\tilde{X}_{i}\}_{i=1}^{n}$ satisfy (26), (27), (28) leading to the convergence in distribution of $\tilde{S}_{n}=\sum_{i}\tilde{X}_{i}$ to $\mathcal{N}\left(0,\sigma^{2}\right)$.
@@
$\begin{split}&\sum_{i=1}^{n}\mathbb{P}\left(|\tilde{X}_{i}|\geq\epsilon\right)=\sum_{i=1}^{n}\left(\int_{-\infty}^{-\epsilon}f_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}+\int_{\epsilon}^{\infty}f_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}\right)=\\ &\sum_{i=1}^{n}\left(\int_{-\infty}^{-c\sqrt{n}\epsilon}f_{X_{i}}\left({x}\right)dx+\int_{c\sqrt{n}\epsilon}^{\infty}f_{X_{i}}\left({x}\right)dx\right)\stackrel{{\scriptstyle n\gg 1}}{{=}}\\ &\sum_{i=1}^{n}\left(\int_{-\infty}^{-c\sqrt{n}\epsilon}\mathcal{O}\left(\frac{1}{x^{4}}\right)dx+\int_{c\sqrt{n}\epsilon}^{\infty}\mathcal{O}\left(\frac{1}{x^{4}}\right)dx\right)=\sum_{i=1}^{n}\left(\mathcal{O}\left(\int_{-\infty}^{-c\sqrt{n}\epsilon}\left|\frac{1}{x^{4}}\right|dx\right)+\mathcal{O}\left(\int_{c\sqrt{n}\epsilon}^{\infty}\left|\frac{1}{x^{4}}\right|dx\right)\right)=\\ &\sum_{i=1}^{n}\mathcal{O}\left(\frac{1}{n^{\frac{3}{2}}}\right)=\mathcal{O}\left(\frac{1}{n^{\frac{1}{2}}}\right)\xrightarrow{n\rightarrow\infty}0.\end{split}$ (32)
@@
$\begin{split}&\sum_{i}^{n}\left(\int_{|\tilde{x}|<\epsilon}\tilde{x}\,f_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}\right)\stackrel{{\scriptstyle n\gg 1}}{{=}}\sum_{i}^{n}\underbrace{\mathbb{E}\left({\tilde{X}_{i}}\right)}_{=0}-\sum_{i}^{n}\left(\mathcal{O}\left(\int_{-\infty}^{-c\sqrt{n}\epsilon}\frac{1}{c\sqrt{n}}\left|\frac{1}{x^{3}}\right|dx\right)+\mathcal{O}\left(\int_{c\sqrt{n}\epsilon}^{\infty}\frac{1}{c\sqrt{n}}\left|\frac{1}{x^{3}}\right|dx\right)\right)=\\ &=\mathcal{O}\left(\frac{1}{n^{\frac{1}{2}}}\right)\xrightarrow{n\rightarrow\infty}0\,.\end{split}$ (33)
@@
$\sum_{i}\left(\int_{|\tilde{x}|<\epsilon}\tilde{x}^{2}\,f_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}-\left(\int_{|\tilde{x}|<\epsilon}\tilde{x}\,f_{\tilde{X}_{i}}\left({\tilde{x}}\right)d\tilde{x}\right)^{2}\right)\stackrel{{\scriptstyle n\gg 1}}{{=}}$ (34)
@@
Then we apply these results to analyze the distribution $f_{Z^{(l+1)}_{i}}^{(\chi)}\left({z{}{}}\right)$; fixed $\mathcal{W}$, in fact, $Z^{(l+1)}_{i}$ can be expressed as a combination of r.v.s. We will demonstrate how the nodes $\left\{Z^{(l+1)}_{i}\right\}_{i}$ may not be identically distributed for $l\geq 1$.
@@ C.2.1 Combination of random variables
To examine the changes in node distribution from one layer to the next, we analyze the distribution of $Z^{(l+1)}_{i}=\sum_{j}w^{(l)}_{ij}H^{(l)}_{j}$, with a fixed set $\mathcal{W}$. This leads to a linear combination of r.v.s. It is straightforward to show that the hypotheses of Thm. C.3 are satisfied in this context. We then have
$f_{Z^{(l+1)}_{i}}^{(\chi)}\left({z{}{}}\right)=\mathcal{N}\left(z{}{};\sum_{j}w^{(l)}_{ij}\left\langle{H^{(l)}_{j}}\right\rangle,\sum_{j}\left(w^{(l)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({H^{(l)}_{j}}\right)\right)$. (39)
Our goal is to understand how $f_{Z^{(l+1)}_{i}}^{(\chi)}\left({z{}{}}\right)$ changes across the set $\left\{Z^{(l+1)}_{i}\right\}_{i=1}^{N_{l+1}}$. Since we are dealing with a Gaussian distribution, it suffices to examine how its mean and variance vary over different nodes $\left\{Z^{(l+1)}_{i}\right\}_{i=1}^{N_{l+1}}$, i.e., using different sets of weights $\left\{w^{(l)}_{i\;\boldsymbol{\cdot}}\right\}_{i=1}^{N_{l+1}}$, with $w^{(l)}_{i\;\boldsymbol{\cdot}}\equiv\left\{w^{(l)}_{ij}\right\}_{j=1}^{N_{l}}$.
Consider the mean $\left\langle{Z^{(l+1)}_{i}}\right\rangle=\sum_{j}w^{(l)}_{ij}\left\langle{H^{(l)}_{j}}\right\rangle$. Now, the set $\left\{\left\langle{H^{(l)}_{j}}\right\rangle\right\}_{j=1}^{N_{l}}$ acts as a fixed constant set, [^10] This is the same set of elements $\forall i\in[1,\dots,N_{l+1}]$. while the set $\left\{w^{(l)}_{i\;\boldsymbol{\cdot}}\right\}_{j=1}^{N_{l}}$ changes with the node index ii. Again, Thm. C.3 is easily seen to be applicable. We get
@@
Assuming $\left\langle{H^{(l)}_{j}}\right\rangle^{2}$ follows the CLT, [^11] The law of large numbers would suffice here as we are not studying convergence to the mean value. we have
$\frac{1}{N_{l}}\sum_{j=1}^{N_{l}}\left\langle{H^{(l)}_{j}}\right\rangle^{2}\xrightarrow{N_{l}\rightarrow\infty}\mathbb{E}_{w^{(l-1)}_{j\boldsymbol{\cdot}}}\left({\left\langle{H^{(l)}_{j}}\right\rangle^{2}}\right)\,$, (41)
@@
$f_{\left\langle{Z^{(l+1)}_{i}}\right\rangle}^{(w^{(l)}_{i\boldsymbol{\cdot}})}\left({z{}{}}\right)\xrightarrow{N_{l}\rightarrow\infty}\mathcal{N}\left(z{}{};0,\sigma_{w}^{2}\mathbb{E}_{w^{(l-1)}_{j\boldsymbol{\cdot}}}\left({\left\langle{H^{(l)}_{j}}\right\rangle^{2}}\right)\right)\,.$ (42)
Regarding the variance in Eq. (39), it will also be a r.v. dependent on the set $\{w^{(l)}_{ij}\}_{j=1}^{N_{l}}$. Following a similar approach to the one used for the sum of Gaussian r.v.s discussed before, we can calculate its mean and variance to show that asymptotically $\sum_{j}\left(w^{(l)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({H^{(l)}_{j}}\right)$ converges to a deterministic value equal for all $N_{l+1}$ nodes.
To make the discussion more concrete we will discuss now, more in detail the specific case of $l=3$ This will be used as an integral part of deriving $f_{G_{0}}\left({g_{0}}\right)$ for deep architectures (see App. G.3.1). In App. G.3.1, we demonstrate that
@@
Our aim is to apply the analysis described above to discern the differences between the set of distributions $\left\{f_{Z^{(3)}_{i}}^{(\chi)}\left({z{}{}}\right)\right\}$. Variations in these distributions are induced by differences in the corresponding sets of weight vectors $\{w^{(2)}_{i\;\boldsymbol{\cdot}}\}$. The sets $\left\{\left\langle{H^{(2)}_{j}}\right\rangle\right\}$ and $\left\{\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)\right\}$, however, remain consistent across each node $Z^{(3)}_{i}$. Therefore, we treat the latter sets as fixed random coefficients, while the variations are attributed to $\{w^{(2)}_{i\;\boldsymbol{\cdot}}\}$. Let us focus on the variance:
$\textrm{Var}_{\chi}\left({Z^{(3)}_{i}}\right)=\sum_{j=1}^{N_{2}}\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)\,$. (44)
To understand how much the random variable $\textrm{Var}_{\chi}\left({Z^{(3)}_{i}}\right)$ varies, we calculate its mean and variance:
$\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\sum_{j=1}^{N_{2}}\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)}\right)=\sigma_{w}^{2}\frac{1}{N_{2}}\sum_{j=1}^{N_{2}}\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)\,\xrightarrow{N_{2}\rightarrow\infty}\sigma_{w}^{2}\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)}\right)\,$, (45)
@@
$\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)}\right)$ $=\int_{\mathbb{R}}\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)(m_{z})\;\;f_{\left\langle{Z^{(2)}_{j}}\right\rangle}^{(w^{(l)}_{i\boldsymbol{\cdot}})}\left({m_{z}}\right)dm_{z}$ (46)
$=\int_{\mathbb{R}}\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)(m_{z})\;\;\mathcal{N}\left(m_{z};0,\sigma_{w}^{2}\mathbb{E}_{w^{(0)}_{j\boldsymbol{\cdot}}}\left({\left\langle{H^{(1)}}\right\rangle^{2}}\right)\right)dm_{z}\,$,
where with $\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)(x)$ we underlined the dependence of $\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)$ from the r.v. $\left\langle{Z^{(2)}_{j}}\right\rangle$. In the last step we substituted the p.d.f. expression from Eq. (42). Note that in the integrand of Eq. (46) there is no scaling dependence with respect to $N_{2}$; this means that
$\mathbb{E}_{w^{(1)}_{j\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)}\right)=\mathcal{O}(1)\Longrightarrow\mathbb{E}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\textrm{Var}_{\chi}\left({Z^{(3)}_{i}}\right)}\right)=\mathcal{O}(1).$ (47)
To evaluate the order of magnitude of the fluctuations, we recall that, given a fixed coefficient $\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)$,
$\textrm{Var}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left(w^{(2)}_{ij}\right)^{2}\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)}\right)=\textrm{Var}_{\chi}\left({H^{(2)}_{j}}\right)^{2}\textrm{Var}_{w^{(2)}_{i\;\boldsymbol{\cdot}}}\left({\left(w^{(2)}_{ij}\right)^{2}}\right)\,.$ (48)